# Nemesis: A Multi-Scale, Multi-Physics Algorithm for Astrophysics

### 03 Mar 2026 v1

@@ Abstract
Aims. In this work, an updated version of the multi-scale, multi-physics algorithm, Nemesis which makes use of the Astrophysical Multipurpose Software Environment (AMUSE). The algorithm is formally introduced and validated.
@@
Conclusions. The method introduced here can find it’s use in numerous domains of astronomy thanks to its flexibility and modularity, from simulating protoplanetary disks in star clusters to binary black holes in the galactic center.
@@ 1 Introduction
In this paper, we formally present a hybrid technique capable in handling multi-physics questions, Nemesis [^1] https://github.com/spzwart/nemesis. This technique was previously used to integrate planetary systems in star clusters (van Elteren et al., 2019), here a more thorough description and validation of the recently modified and optimised version is provided in detail. Results show that the core approximation on which Nemesis is built, that is, decoupling widely separated systems from each other and solving them in parallel, yields negligible loss in physics at the reward of massive computational benefits. The flexibility this hybrid code provides to the user allows researchers to tackle multi-scale and multi-physics problems.
@@ 2.1 Dealing with Scales
Over time, the number of children, $N_{\rm chd}$, may vary. During the parent integration, close encounters form a new child system. Conversely, a child system dissolves if any of its members become sufficiently separated (i.e an ionising event between a binary and an intruder occurs). If $N_{\rm chd}$ exceeds the number of cores available, excess children wait in a queue until a CPU becomes available. CPUs are released the moment a child code finishes integrating, with priority given to the most computationally expensive jobs.
@@ 2.2 Synchronising Scales
    While the corrections are minor, they accumulate in time. Moreover, as mentioned earlier, large NN systems are inherently chaotic (Poincaré, 1891, 1892; Dejonghe and Hut, 1986). As such, inclusion of these corrections may translate into subtle, yet distinct dynamical effects.
@@
    As before, to give context with numbers, a virialised Plummer sphere of density $100$ M pc-3 having $10/200$ of the stars hosting planetary systems, the corrections applied to the planets over the same $500$ yr bridge time steps results in velocity shifts of $\delta v\sim 10^{-10}-10^{-7}$ km s-1. The large spread is due to the sensitivity of the local environment with gravity having an inverse-squared law with spatial separations. When $\rho\sim 10^{3}$ M pc-3, this ranges between $\delta v\sim 10^{-8}-10^{-5}$ km s-1.
Correction kicks are applied every bridge time step via a second-order Verlet kick-drift-kick method (Verlet, 1967; Fujii et al., 2007). In the limit $\delta t_{\rm nem}\xrightarrow[]{}0$ yr, Nemesis mimics a direct 22nd-Order NN-body integrator, although one with substantial overhead. In the opposite limit, $\delta t_{\rm nem}\xrightarrow$$\infty$ yr, computational time is minimised but accuracy is lost.
@@
Once a particle is no longer identified as a child, for example when it is ejected and becomes a rogue planet, it becomes a parent, after which its dynamical evolution will be treated in the parent code and it will feel the effects of all still existing children through correction kicks (see section 2.4).
@@ 2.3 Handling Parent Mergers
Here $A$ is a user-defined coefficient (default $A=100$) and $M_{\rm par}$ is the mass of the parent. For isolated parents (i.e., parents hosting no children, such as a rogue planet or single star) $M_{\rm par}$ is the particle’s mass. If, instead, the parent hosts children, $M_{\rm par}$ is the total mass of the child system. By default, an upper limit of $R_{\rm par}=10^{3}$ au is applied. While test particles are massless and therefore have no radii, isolated test particles can still merge upon the approach of another (massive) parent particle. Nemesis handles mergers using two general schemes:
@@
  • Case 2: Neither parent particles host children
@@ 2.4 Handling Children Dissolutions
Just as parent particles can approach one another and form a new child system, child particles can also move apart. This can happen, for instance, with high velocity stars in the galactic centre or ejected debris in protoplanetary disks etc…
@@
At each bridge time step, particles are grouped into connected components by linking all pairs with separations $r_{ij}\leq\alpha R_{\rm par}$. The set of particles attributed to a connected component forms a child system, and will have its internal dynamics resolved using a dedicated child code. If, during the evolution, a child system fragments into multiple connected components, the system is flagged as dissolved and is handled as follows:
@@ 2.5 Code Coupling
Stellar evolution defaults to using the parameterised code SeBa (Portegies Zwart and Verbunt, 1996; Portegies Zwart and Yungelson, 1998; Nelemans et al., 2001; Toonen et al., 2012). SeBa can handle binary evolution and provides a quick way to resolve the general evolution of a star, although it doesn’t provide its internal structure. A stopping condition is enabled to flag any star undergoing a supernova. If triggered, a natal kick is applied to the remnant star in focus. If one wishes to reproduce the rejuvenation and collision of stars better, one can adopt a Henyey solver (i.e MESA, Paxton et al. (2011)) instead. Figure 1 shows the general workflow of Nemesis, a video can also be accessed online [^2] https://youtu.be/cycIn8hDZKY.
Figure 1: Nemesis workflow. star_evol is a parameter in Nemesis toggling on or off stellar evolution. Communication between codes is internally handled by AMUSE via channels. Orange boxes represent the setup and termination of the simulation, red boxes indicate steps where work is offloaded to external codes, and blue boxes denote tasks handed back to Python for manipulation of Nemesis-structured data.
@@ 3.3 Computational Scaling: Nemesis Scaling Relations
The last set of runs investigates the computational scaling relations of Nemesis. The simulated cluster environment is set up similarly to that in section 3.1. Planets follow the oligarchic model once more, however, the number is increased to eight per system, and asteroids are no longer present. Three cluster configurations are considered: A massive, dense cluster with $N_{*}=2048$ and $r_{h}=0.5$ pc, a low-populated, sparse cluster with $N_{*}=256$ and $r_{h}=0.5$ pc and, finally, a massive, sparse cluster with $N_{*}=2048$ and $r_{h}=1$ pc. The latter two have equal densities. In all cases, the cluster is sampled from a Plummer distribution and initially in virial equilibrium. Clusters are integrated until $t=0.1$ Myr.
When analysing the dependence of computational resources on $\delta t_{\rm Nem}$, the number of planetary systems is initially fixed to $N_{\rm chd}=20$ in all three configurations while $\delta t_{\rm nem}$ varies between $10^{2}-10^{3}$ yrs, in equal increments of $10^{2}$ yr. When testing how the number of children influences computational resources, only the massive, dense cluster is considered and the bridge time step is fixed to $\delta t_{\rm Nem}=10^{3}$ yr. In this case, $N_{\rm chd}$ initially ranges between 206020-60, spaced out in equal increments of ten. Although the same cluster realisation is used for the initial condition in all runs for the latter test, by changing $N_{\rm chd}$, some runs have stars hosting planetary systems which otherwise do not in others.
@@ 4 Results and Discussion
Figure 2: Evolution of simulation’s energy error in time.
@@
The advantages provided by Nemesis’s flexibility in allowing for a symplectic algorithm for its set of children is evident in figure 2 since its relative energy error remains essentially bounded (see figure A.4 of van Elteren et al. (2019) for another example). Jumps, however, still appear due to the symplectic nature of Huayno hindering its ability to resolve close approaches.
@@
A Cramer-von Misses (Cramér, 1928) test is done to see whether there are statistically important differences in results. This test can compare the distribution spanning the total range of values, something lacking in the Kolmogorov-Smirnov test which only probes the region with largest deviation. A $p$-value of $p_{a}=0.98$ and $p_{e}=0.93$ is found for $a$ and $e$, respectively, implying that the distributions are indistinguishable.
@@
Given the importance of the tail-end distribution in an astronomical context since high-$e$ orbits can induce system instabilities, and/or result in mergers/ejection events, a two-sample Anderson and Darling (1954) test is also used. In the end, a $p$-value of $p_{a}=0.93$ and $p_{e}=0.50$ is found, allowing us to confidently reject the null-hypothesis that either samples are drawn from different distributions. The large-scale cluster properties also exhibits similar results with both the distribution in velocities and position of stars relative to one another having $p=1$.
@@
The residuals remain low and are likely attributed to statistical noise, a feature which naturally arises given the chaotic nature of many-body dynamical systems. Regardless, it is necessary to mention that Nemesis does perform worst when resolving the dynamical evolution of particles with semi-major axis $a\sim R_{\rm par}$, as indicated by the top panel of figure 4. The origin of this discrepancy lies in the synchronisation between children and the global environment occurring at discrete time intervals (recall section 2.2).
In Nemesis, correction kicks are only applied once every fixed bridge time-steps, $\delta t_{\rm nem}$. These bridge time steps are significantly larger than the internal integration time-step $\delta t$ used by gravitational solvers such as Ph4 (equation 4). Indeed, while Ph4 accounts for the gravitational influence of all particles at every internal time-step, child particles in Nemesis experience the external potential only at the bridge time steps. In between these bridge time steps, child particles only interact with those part of the same child system.
By having poorer resolution of external interactions, and being more likely to compute the potential when fly-by events are already relatively far away, Nemesis underestimates the effective potential. Naturally, it is the child particles orbiting at $a\sim R_{\rm par}$ which are the most affected by this reduced resolution of the external potential.
@@
A direct $N$-body code can accurately resolve this encounter by integrating the equations of motion at each internal time-step (represented schematically by tracks on the fly-by trajectory). In contrast, under the fixed bridge implementation of Nemesis where each correction kick is computed at increments of $\delta t_{\rm nem}$, close interactions are likely to be under-resolved or missed. Consequently, although correction kicks are applied, the algorithm predominantly samples the external potential when the perturber is already relatively distant. This effect is only significant in moderate encounters. In the case that $\delta r_{ij}\leq R_{{\rm par},i}+R_{{\rm par},j}$, the two systems will merge and all relevant particles will form a new children system where their internal dynamics will be resolved using a dedicated symplectic integrator (section 2.3).
@@
Figure 6 shows the oscillation of the total inclination ($i_{\rm tot}=i_{\rm out}+i_{\rm in}$) and the inner binaries eccentricity over time for all three runs. Runs using a test particle as the planetary mass object yield identical results and in both cases, Nemesis resolves the ZLK effect over secular times at an accuracy indistinguishable from a direct $N$-body integrator (the inlet in the lower panel highlights the extent to which Nemesis replicates results from a direct $N$-body integration). A relative energy error of $\Delta E=3\times 10^{-12}$ is found at the final step.
Theory estimates that the system considered here has an ZLK oscillation period of $0.913$ Myr (Naoz et al., 2013b). The estimate uses a double-averaged Hamiltonian, and as such is an approximation. Additionally, they note that when the inner binary reaches large eccentricities (as the case here), the approximation can change by ‘orders-of-magnitudes’. Even so, the value is near the oscillation timescale seen here ($0.821$ Myr). Even with the consistent splitting of the child system and merging of parent particles, the procedure doesn’t affect the dynamics, with a smooth evolution in the orbital parameters being observed. The result also highlights the validity of representing children with a CoM parent in the parent code.
@@ 4.3.1 Bridge Time Step
There is no optimum time step. The user should conduct test runs to find the optimal parameter. After all, a balancing act between computational resources saved by increasing $\delta t_{\rm nem}$ and loss of accuracy exists. Ideally, Nemesis would use an adaptive algorithm. For instance, as a cluster evolves and expands, long-range perturbations become increasingly weaker such that $\delta t_{\rm nem}$ can relax to larger values. As mentioned in the end of section 4.1, although an adaptive time step scheme is not yet implemented, it is in the plans for future work.
@@ 4.3.2 Number of Children
Table 1: Wall-clock time taken to simulate $N_{\rm chd}$ systems till $0.1$ Myr. Col. 1: The number of children. Col. 2: Results using the direct $N$-body algorithm. Col. 3: Results using Nemesis.
@@ 4.4 Parent Radius
    Avoiding Bottlenecks from Tight Orbits: The problem of directly integrating small-scale systems in large-scale environments (i.e planetary systems in clusters) is the need to suppress the drift in energy errors to allow capture of secular effects. This can be done with symplectic codes. However, with their shared time step scheme, the tightest orbits induce a bottleneck in the integration. With Nemesis, as long as $R_{\rm par}$ exceeds a typical system’s tightest orbit, tight orbits are always decoupled from the global integration, removing the bottleneck for the parent integration while allowing for accurate integration of the internal dynamics. This scheme is indifferent to the extent $R_{\rm par}$ exceeds the scale of the tightest orbit.
@@
Throughout testing, it was found that a radius coefficient $A=100$ sufficed (see equation 3). This was tested for environments ranging in densities between $\rho\sim(10^{2}-10^{4})\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ but may change depending on the initial cluster parameters, i.e initial virality or compactness of systems.
@@
A conservative $R_{\rm par}$ implies that wide systems, i.e., wide planets or Oort Cloud-like environments (Öpik, 1932; Oort, 1950), fall back to the parent code. While this does not resolve the local dynamics as successfully, the long secular timescales of wide orbits partly compensate for the loss in accuracy. Moreover, such systems are inherently fragile, and external tides or encounters with field particles are likely to strip them before secular effects become significant. There is a delicate balance between $\delta t_{\rm nem}$ and $R_{\rm par}$, and it remains for future work to find a logical and adaptable way to balance the two for a given system keeping in mind computation time and accuracy of results.

### 10 Mar 2026 v2

@@ Abstract
Aims. In this work, an updated version of the multi-scale, multi-physics algorithm, Nemesis, which makes use of the Astrophysical Multipurpose Software Environment (AMUSE). The algorithm is formally introduced and validated.
@@
Conclusions. The method introduced here can find its use in numerous domains of astronomy, thanks to its flexibility and modularity, from simulating protoplanetary disks in star clusters to binary black holes in the galactic centre.
@@ 1 Introduction
In this paper, we formally present a hybrid technique capable of handling multi-physics questions, Nemesis [^1] https://github.com/spzwart/nemesis. This technique was previously used to integrate planetary systems in star clusters (van Elteren et al., 2019); here, a more thorough description and validation of the recently modified and optimised version is provided in detail. Results show that the core approximation on which Nemesis is built, that is, decoupling widely separated systems from each other and solving them in parallel, yields negligible loss in physics at the reward of massive computational benefits. The flexibility this hybrid code provides to the user allows researchers to tackle multi-scale and multi-physics problems.
@@ 2.1 Dealing with Scales
Over time, the number of children, $N_{\rm chd}$, may vary. During the parent integration, close encounters form a new child system. Conversely, a child system dissolves if any of its members become sufficiently separated (i.e., an ionising event between a binary and an intruder occurs). If $N_{\rm chd}$ exceeds the number of cores available, excess children wait in a queue until a CPU becomes available. CPUs are released the moment a child code finishes integrating, with priority given to the most computationally expensive jobs.
@@ 2.2 Synchronising Scales
    While the corrections are minor, they accumulate in time. Moreover, as mentioned earlier, large NN systems are inherently chaotic (Poincaré, 1891, 1892; Dejonghe and Hut, 1986). As such, inclusion of these corrections may translate into subtle yet distinct dynamical effects.
@@
    As before, to give context with numbers, a virialised Plummer sphere of density $100$ M pc-3 having $10/200$ of the stars hosting planetary systems, the corrections applied to the planets over the same $500$ yr bridge time steps results in velocity shifts of $\delta v\sim 10^{-10}-10^{-7}$ km s-1. The large spread is due to the sensitivity of the local environment, with gravity having an inverse-squared law with spatial separations. When $\rho\sim 10^{3}$ M pc-3, this ranges between $\delta v\sim 10^{-8}-10^{-5}$ km s-1.
Correction kicks are applied every bridge time step via a second-order Verlet kick-drift-kick method (Verlet, 1967; Fujii et al., 2007). In the limit $\delta t_{\rm nem}\xrightarrow[]{}0$ yr, Nemesis mimics a direct 22nd-Order NN-body integrator, although one with substantial overhead. In the opposite limit, $\delta t_{\rm nem}\xrightarrow$$\infty$ yr, computational time is minimised, but accuracy is lost.
@@
Once a particle is no longer identified as a child, for example, when it is ejected and becomes a rogue planet, it becomes a parent, after which its dynamical evolution will be treated in the parent code, and it will feel the effects of all still existing children through correction kicks (see section 2.4).
@@ 2.3 Handling Parent Mergers
Here $A$ is a user-defined coefficient (default $A=100$) and $M_{\rm par}$ is the mass of the parent. For isolated parents (i.e., parents hosting no children, such as a rogue planet or single star), $M_{\rm par}$ is the particle’s mass. If, instead, the parent hosts children, $M_{\rm par}$ is the total mass of the child system. By default, an upper limit of $R_{\rm par}=10^{3}$ au is applied. While test particles are massless and therefore have no radii, isolated test particles can still merge upon the approach of another (massive) parent particle. Nemesis handles mergers using two general schemes:
@@
  • Case 2: Neither parent particle hosts children
@@ 2.4 Handling Children Dissolutions
Just as parent particles can approach one another and form a new child system, child particles can also move apart. This can happen, for instance, with high velocity stars in the galactic centre or ejected debris in protoplanetary disks, etc…
@@
At each bridge time step, particles are grouped into connected components by linking all pairs with separations $r_{ij}\leq\alpha R_{\rm par}$. The set of particles attributed to a connected component forms a child system and will have its internal dynamics resolved using a dedicated child code. If, during the evolution, a child system fragments into multiple connected components, the system is flagged as dissolved and is handled as follows:
@@ 2.5 Code Coupling
Stellar evolution defaults to using the parameterised code SeBa (Portegies Zwart and Verbunt, 1996; Portegies Zwart and Yungelson, 1998; Nelemans et al., 2001; Toonen et al., 2012). SeBa can handle binary evolution and provides a quick way to resolve the general evolution of a star, although it doesn’t provide its internal structure. A stopping condition is enabled to flag any star undergoing a supernova. If triggered, a natal kick is applied to the remnant star in focus. If one wishes to reproduce the rejuvenation and collision of stars better, one can adopt a Henyey solver (i.e MESA, Paxton et al. (2011)) instead. Figure 1 shows the general workflow of Nemesis. A video can also be accessed online.
Figure 1: Nemesis workflow. star_evol is a parameter in Nemesis toggling on or off stellar evolution. Communication between codes is internally handled by AMUSE via channels. Orange boxes represent the setup and termination of the simulation, red boxes indicate steps where work is offloaded to external codes, and blue boxes denote tasks handed back to Python for manipulation of Nemesis-structured data. A movie showing Nemesis in action is available online.
@@ 3.3 Computational Scaling: Nemesis Scaling Relations
The last set of runs investigates the computational scaling relations of Nemesis. The simulated cluster environment is set up similarly to that in section 3.1. Planets follow the oligarchic model once more; however, the number is increased to eight per system, and asteroids are no longer present. Three cluster configurations are considered: A massive, dense cluster with $N_{*}=2048$ and $r_{h}=0.5$ pc, a low-populated, sparse cluster with $N_{*}=256$ and $r_{h}=0.5$ pc and, finally, a massive, sparse cluster with $N_{*}=2048$ and $r_{h}=1$ pc. The latter two have equal densities. In all cases, the cluster is sampled from a Plummer distribution and initially in virial equilibrium. Clusters are integrated until $t=0.1$ Myr.
When analysing the dependence of computational resources on $\delta t_{\rm Nem}$, the number of planetary systems is initially fixed to $N_{\rm chd}=20$ in all three configurations while $\delta t_{\rm nem}$ varies between $10^{2}-10^{3}$ yrs, in equal increments of $10^{2}$ yr. When testing how the number of children influences computational resources, only the massive, dense cluster is considered, and the bridge time step is fixed to $\delta t_{\rm Nem}=10^{3}$ yr. In this case, $N_{\rm chd}$ initially ranges between 206020-60, spaced out in equal increments of ten. Although the same cluster realisation is used for the initial condition in all runs for the latter test, by changing $N_{\rm chd}$, some runs have stars hosting planetary systems which otherwise do not in others.
@@ 4 Results and Discussion
Figure 2: Evolution of the simulation’s energy error in time.
@@
The advantages provided by Nemesis’s flexibility in allowing for a symplectic algorithm for its set of children are evident in figure 2 since its relative energy error remains essentially bounded (see figure A.4 of van Elteren et al. (2019) for another example). Jumps, however, still appear due to the symplectic nature of Huayno hindering its ability to resolve close approaches.
@@
A Cramer-von Misses (Cramér, 1928) test is done to see whether there are statistically significant differences in results. This test can compare the distribution spanning the total range of values, something lacking in the Kolmogorov-Smirnov test, which only probes the region with the largest deviation. A $p$-value of $p_{a}=0.98$ and $p_{e}=0.93$ is found for $a$ and $e$, respectively, implying that the distributions are indistinguishable.
@@
Given the importance of the tail-end distribution in an astronomical context since high-$e$ orbits can induce system instabilities, and/or result in mergers/ejection events, a two-sample Anderson and Darling (1954) test is also used. In the end, a $p$-value of $p_{a}=0.93$ and $p_{e}=0.50$ is found, allowing us to confidently reject the null-hypothesis that either samples are drawn from different distributions. The large-scale cluster properties also exhibit similar results with both the distribution in velocities and position of stars relative to one another having $p=1$.
@@
The residuals remain low and are likely attributed to statistical noise, a feature which naturally arises given the chaotic nature of many-body dynamical systems. Regardless, it is necessary to mention that Nemesis does perform worse when resolving the dynamical evolution of particles with semi-major axis $a\sim R_{\rm par}$, as indicated by the top panel of figure 4. The origin of this discrepancy lies in the synchronisation between children and the global environment occurring at discrete time intervals (recall section 2.2).
In Nemesis, correction kicks are only applied once every fixed bridge time-steps, $\delta t_{\rm nem}$. These bridge time steps are significantly larger than the internal integration time step $\delta t$ used by gravitational solvers such as Ph4 (equation 4). Indeed, while Ph4 accounts for the gravitational influence of all particles at every internal time step, child particles in Nemesis experience the external potential only at the bridge time steps. In between these bridge time steps, child particles only interact with those part of the same child system.
By having poorer resolution of external interactions and being more likely to compute the potential when fly-by events are already relatively far away, Nemesis underestimates the effective potential. Naturally, it is the child particles orbiting at $a\sim R_{\rm par}$ which are the most affected by this reduced resolution of the external potential.
@@
A direct $N$-body code can accurately resolve this encounter by integrating the equations of motion at each internal time step (represented schematically by tracks on the fly-by trajectory). In contrast, under the fixed bridge implementation of Nemesis, where each correction kick is computed at increments of $\delta t_{\rm nem}$, close interactions are likely to be under-resolved or missed. Consequently, although correction kicks are applied, the algorithm predominantly samples the external potential when the perturber is already relatively distant. This effect is only significant in moderate encounters. In the case that $\delta r_{ij}\leq R_{{\rm par},i}+R_{{\rm par},j}$, the two systems will merge and all relevant particles will form a new children system where their internal dynamics will be resolved using a dedicated symplectic integrator (section 2.3).
@@
Figure 6 shows the oscillation of the total inclination ($i_{\rm tot}=i_{\rm out}+i_{\rm in}$) and the inner binaries’ eccentricity over time for all three runs. Runs using a test particle as the planetary mass object yield identical results and in both cases, Nemesis resolves the ZLK effect over secular times at an accuracy indistinguishable from a direct $N$-body integrator (the inlet in the lower panel highlights the extent to which Nemesis replicates results from a direct $N$-body integration). A relative energy error of $\Delta E=3\times 10^{-12}$ is found at the final step.
Theory estimates that the system considered here has a ZLK oscillation period of $0.913$ Myr (Naoz et al., 2013b). The estimate uses a double-averaged Hamiltonian, and as such is an approximation. Additionally, they note that when the inner binary reaches large eccentricities (as the case here), the approximation can change by ‘orders of magnitude’. Even so, the value is near the oscillation timescale seen here ($0.821$ Myr). Even with the consistent splitting of the child system and merging of parent particles, the procedure doesn’t affect the dynamics, with a smooth evolution in the orbital parameters being observed. The result also highlights the validity of representing children with a CoM parent in the parent code.
@@ 4.3.1 Bridge Time Step
There is no optimum time step. The user should conduct test runs to find the optimal parameter. After all, a balancing act between computational resources saved by increasing $\delta t_{\rm nem}$ and loss of accuracy exists. Ideally, Nemesis would use an adaptive algorithm. For instance, as a cluster evolves and expands, long-range perturbations become increasingly weaker such that $\delta t_{\rm nem}$ can relax to larger values. As mentioned at the end of section 4.1, although an adaptive time step scheme is not yet implemented, it is in the plans for future work.
@@ 4.3.2 Number of Children
Table 1: Wall-clock time taken to simulate $N_{\rm chd}$ systems till $0.1$ Myr. [^2] Col. 1: The number of children. Col. 2: Results using the direct $N$-body algorithm. Col. 3: Results using Nemesis.
@@ 4.4 Parent Radius
    Avoiding Bottlenecks from Tight Orbits: The problem of directly integrating small-scale systems in large-scale environments (i.e., planetary systems in clusters) is the need to suppress the drift in energy errors to allow capture of secular effects. This can be done with symplectic codes. However, with their shared time step scheme, the tightest orbits induce a bottleneck in the integration. With Nemesis, as long as $R_{\rm par}$ exceeds a typical system’s tightest orbit, tight orbits are always decoupled from the global integration, removing the bottleneck for the parent integration while allowing for accurate integration of the internal dynamics. This scheme is indifferent to the extent $R_{\rm par}$ exceeds the scale of the tightest orbit.
@@
Throughout testing, it was found that a radius coefficient $A=100$ sufficed (see equation 3). This was tested for environments ranging in densities between $\rho\sim(10^{2}-10^{4})\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ but may change depending on the initial cluster parameters, i.e., initial virality or compactness of systems.
@@
A conservative $R_{\rm par}$ implies that wide systems, i.e., wide planets or Oort Cloud-like environments (Öpik, 1932; Oort, 1950), fall back to the parent code. While this does not resolve the local dynamics as successfully, the long secular timescales of wide orbits partly compensate for the loss in accuracy. Moreover, such systems are inherently fragile, and external tides or encounters with field particles are likely to strip them before secular effects become significant. There is a delicate balance between $\delta t_{\rm nem}$ and $R_{\rm par}$, and it remains for future work to find a logical and adaptable way to balance the two for a given system, keeping in mind computation time and accuracy of results.